# Extending Linear Convergence of the Proximal Point Algorithm: The Quasar-Convex Case

### v1

@@ 1 Introduction
Despite its simplicity, PPA is a powerful tool in continuous optimization, as it allows us to simplify the original problem by solving better-conditioned subproblems. For instance, when $h$ is convex, the function $x\mapsto h(x)+\frac{1}{2\beta_{k}}\lVert x^{k}-x\rVert^{2}$ is strongly convex. This property makes solving the subproblems easier because proper, lower semicontinuous (lsc henceforth) strongly convex functions possess remarkable properties, such as a unique minimizer and quadratic growth. In fact, when $h$ is convex, the sequence $\{x^{k}\}_{k}$ generated by (2) is a minimizing sequence for problem (1). Specifically, it is decreasing, converges to the minimum (though not linearly), and achieves a complexity rate of $\mathcal{O}(\varepsilon^{-1})$. Moreover, if $h$ is strongly convex, then $\{x^{k}\}_{k}$ converges linearly to the unique solution with an improved complexity rate of $\mathcal{O}(\ln(\varepsilon^{-1}))$.
@@
Our Contribution: We study (strongly) quasar-convex functions with modulus $\gamma\geq 0$. To that end, first we establish sufficient conditions for (nonsmooth) functions to be (strongly) quasar-convex. In particular, we show that the $\ell_{p}$-regularization ($0<p<1$) [52] and the CES (Constant Elasticity of Substitution) [3], Cobb-Douglas [12], and Leontief [33] production/utility functions are quasar-convex (and strongly quasar-convex on bounded convex sets). Second, we investigate the proximity operator of quasar-convex functions by proving that it is always nonempty (i.e., existence of iterates is guaranteed) and demonstrating cases where it is not a singleton. Furthermore, we establish a useful relationship between the minimizers of the proximity operator and the minimizers of the function. Third, we prove that the set of fixed points of the proximity operator coincides with the set of minimizers of a quasar-convex function, which provides the stopping criterion for PPA. Fourth, we implement PPA for quasar-convex functions ($\gamma\geq 0$) and show that the generated sequence is minimizing with a complexity rate of $\mathcal{O}(\varepsilon^{-1})$ when $\gamma=0$, matching the rate for convex functions [19]. Fifth, for the special case of strongly quasar-convex functions ($\gamma>0$), we prove that PPA converges linearly to the unique solution and achieves an $\mathcal{O}(\ln(\varepsilon^{-1}))$ complexity rate, exactly as in the strongly convex case. Finally, we present numerical experiments with nonconvex examples to demonstrate the potential of our results.
@@ 2 Preliminaries and Basic Definitions
The inner product of $\mathbb{R}^{n}$ and the Euclidean norm are denoted by $\langle\cdot,\cdot\rangle$ and $\lVert\cdot\rVert$, respectively. The open ball with center at $x_{0}\in\mathbb{R}^{n}$ and radius $\varepsilon>0$ is denoted by $\mathbb{B}(x_{0},\varepsilon)$. We denote $\mathbb{R}_{+}:=[0,+\infty[$ and $\mathbb{R}_{++}:=\,]0,+\infty[$, hence $\mathbb{R}^{n}_{+}:=[0,+\infty[\,\times\ldots\times[0,+\infty[$ and $\mathbb{R}_{++}^{n}:=\,]0,+\infty[\,\times\ldots\times\,]0,+\infty[$ ($n$ times), respectively. Let $K$ be a nonempty set in $\mathbb{R}^{n}$, its closure is denoted by $\overline{K}$ and its asymptotic (recession) cone by $K^{\infty}:=\{u\in\mathbb{R}^{n}:~\exists~t_{k}\rightarrow+\infty,~\exists~x_{k}\in K,~\frac{x_{k}}{t_{k}}\rightarrow u\}$. Furthermore, for any set $K\subset\mathbb{R}^{n}$, it follows from [5, Proposition 2.1.2] that $K^{\infty}=\{0\}$ if and only if $K$ is bounded.
@@
    (strongly) quasiconvex on ${\rm dom}\,h$, if there exists $\gamma\geq 0$ such that if for all $x,y\in{\rm dom}\,h$ and all $\lambda\in[0,1]$, we have
@@
A proper function $h:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is said to be:
@@
    2-supercoercive, if
    $\liminf_{\lVert x\rVert\rightarrow+\infty}\frac{h(x)}{\lVert x\rVert^{2}}>0,$ (7)
@@
    coercive, if
    $\lim_{\lVert x\rVert\rightarrow+\infty}h(x)=+\infty.$ (8)
    or equivalently, if $S_{\lambda}(h)$ is bounded for all $\lambda\in\mathbb{R}$.
@@
The following generalized convexity notion has been used in several applications from machine learning in virtue of its properties for fast convergence of gradient-type method (see [20, 23, 24, 25, 50] and references therein).
@@
Let $\kappa\in\,]0,1]$, $\gamma\geq 0$ and $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a function. Then $h$ is said to be $(\kappa,\gamma)$-strongly quasar-convex with respect to $\overline{x}\in{\rm argmin}_{\mathbb{R}^{n}}\,h$ if
$h(\lambda\overline{x}+(1-\lambda)x)\leq\kappa\lambda h(\overline{x})+(1-\kappa\lambda)h(x)-\lambda\left(1-\frac{\lambda}{2-\kappa}\right)\frac{\kappa\gamma}{2}\lVert x-\overline{x}\rVert^{2},$
for all $\lambda\in[0,1]$ and all $x\in\mathbb{R}^{n}$. Furthermore, when $\gamma=0$, we simply say that $h$ is $\kappa$-quasar-convex with respect to $\overline{x}\in{\rm argmin}_{\mathbb{R}^{n}}\,h$.
Note that ${\rm argmin}_{\mathbb{R}^{n}}\,h$ is a singleton when $h$ is strongly quasar-convex ($\gamma>0$), thus in this case we say that $h$ is strongly quasar-convex with modulus $\kappa\in\,]0,1]$ and $\gamma>0$. If $h$ is quasar-convex ($\gamma=0$), then ${\rm argmin}_{\mathbb{R}^{n}}\,h$ is star-shaped (see [25, Observation 3]), i.e., there exists an $\overline{x}_{0}\in{\rm argmin}_{\mathbb{R}^{n}}\,h$ such that for all $\overline{x}\in{\rm argmin}_{\mathbb{R}^{n}}\,h$ and all $\lambda\in[0,1]$, we have $\lambda\overline{x}_{0}+(1-\lambda)\overline{x}\in{\rm argmin}_{\mathbb{R}^{n}}\,h$. Furthermore, if $\kappa=1$, then (strongly) quasar-convex functions are (strongly) star-convex [42]. In particular, every (strongly) convex function is $(1,\gamma)$-strongly quasar-convex ($\gamma\geq 0$).
@@
where “*” denotes that ${\rm argmin}_{\mathbb{R}^{n}\,}h\neq\emptyset$ is required. All the reverse statements do not hold in general (see [20, 24, 25, 42]).
@@
Note that if $h$ is strongly quasar-convex with modulus $\kappa\in\,]0,1]$ and $\gamma>0$, then (see [25, Corollary 1])
$h(\overline{x})+\frac{\kappa\gamma}{2(2-\kappa)}\lVert y-\overline{x}\rVert^{2}\leq h(y),~\forall~y\in\mathbb{R}^{n},$ (11)
where $\overline{x}\in{\rm argmin}_{\mathbb{R}^{n}}\,h$, thus strongly quasar-convex functions satisfy a quadratic growth condition with modulus $\frac{\kappa\gamma}{2(2-\kappa)}>0$.
@@ 3.1 Properties and Examples
We begin this subsection with the following simple results.
@@
    The function $h$ is $\kappa$-quasar-convex with respect to $\overline{x}=0$ if and only if
@@
    The function $h$ is $(\kappa,\gamma)$-strongly quasar-convex with respect to $\overline{x}=0$ if and only if there exist $\kappa\in\,]0,1]$ and $\gamma>0$ such that (16) holds and
    $0<\gamma\leq Q_{\alpha,\kappa}(\lambda)\frac{h(x)}{\lVert x\rVert^{2}},~\forall~x\in K\setminus\{0\},~\forall~\lambda\in\,]0,1].$ (17)
@@
Since $\overline{x}=0\in{\rm argmin}_{K}\,h$ and $h$ satisfies (15), $h(0)=0$. Thus, $h$ is $(\kappa,\gamma)$-strongly quasar-convex with respect to $\overline{x}=0$ if and only if there exist $\kappa\in\,]0,1]$ and $\gamma>0$ with
@@
    If $0<\alpha<1$, then $h$ is $\kappa$-quasar-convex with respect to $\overline{x}=0$ for every $\kappa\in\,]0,\alpha]$.
@@
    If $\alpha\geq 1$, then $h$ is $\kappa$-quasar-convex with respect to $\overline{x}=0$ for every $\kappa\in\,]0,1]$.
@@
    If $0<\alpha\leq 1$, then $h$ is $(\kappa,\gamma)$-strongly quasar-convex with respect to $\overline{x}=0$ for every $\kappa\in\,]0,\alpha[$ and
@@
    If $1<\alpha<2$, then $h$ is $(\kappa,\gamma)$-strongly quasar-convex with respect to $\overline{x}=0$ for every $\kappa\in\,]0,1]$ and
@@
    If $\alpha=2$, then $h$ is $(\kappa,\gamma)$-strongly quasar-convex with respect to $\overline{x}=0$ for every $\kappa\in\,]0,1]$ and
@@
Since $K_{c}:=\{x\in K:\lVert x\rVert=c\}$ is compact and $h$ is lsc, there exists $x_{0}\in K_{c}$ such that
$h(x_{0})=\min_{x\in K_{c}}h(x)=S_{c}.$
As $x_{0}\neq 0$ and $\mathrm{argmin}_{K}\,h=\{0\}$, it follows that $S_{c}>0$. Let $x\in K\setminus\{0\}$ and take $v:=\frac{cx}{\lVert x\rVert}$. By assumption (19), we have $v\in K$ whenever $\lVert x\rVert\leq c$, thus the convexity of $K$ ensures $v\in K$ when $\lVert x\rVert>c$. Then we have two cases:
@@
$h_{p}(x)=\lVert x\rVert_{p}=\left(\sum^{n}_{i=1}\lvert x_{i}\rvert^{p}\right)^{\frac{1}{p}}.$ (20)
@@
It follows from Proposition 5 that $h_{p}$ is quasar-convex on $\mathbb{R}^{n}$. In particular, when $K=\overline{\mathbb{B}}(0,1)$ and $c=1$, we have $S_{1}=\displaystyle\min_{\lVert x\rVert=1}~\lVert x\rVert_{p}=1$. Hence, the modulus $\gamma$ and $\kappa$ given by Proposition 6 are
@@ 3.2 On the Proximity Operator
The next result provide an analysis of the behaviour at infinity of strongly quasar-convex functions.
@@
Let $K\subset\mathbb{R}^{n}$ be a closed convex set and $h:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ be a proper lsc function such that $K\subset{\rm dom}\,h$. If $h$ is a strongly quasar-convex function with modulus $\kappa\in\,]0,1]$ and $\gamma>0$, then $h$ is $2$-supercoercive.
@@
Let $\{x_{k}\}_{k}\subset K$ be such that $\lVert x_{k}\rVert\rightarrow+\infty$. Then, we may assume that $\frac{x_{k}}{\lVert x_{k}\rVert}\rightarrow u\in K^{\infty}$ with $\lVert u\rVert=1$. Since $h$ has a minimum $\overline{x}\in K$, $h(\overline{x})\leq h(x_{k})$ for all $k\in\mathbb{N}$. Using relation (11), we have
$h(x^{k})\geq h(\overline{x})+\frac{\kappa\gamma}{2(2-\kappa)}\lVert x_{k}-\overline{x}\rVert^{2}$ $\Longrightarrow\frac{h(x_{k})}{\lVert x_{k}\rVert^{2}}\geq\frac{h(\overline{x})}{\lVert x_{k}\rVert^{2}}+\frac{\kappa\gamma}{2(2-\kappa)}\left\|\frac{x_{k}}{\lVert x_{k}\rVert}-\frac{\overline{x}}{\lVert x_{k}\rVert}\right\|^{2}$
$\Longrightarrow\liminf_{k\rightarrow+\infty}\frac{h(x_{k})}{\lVert x_{k}\rVert^{2}}\geq\frac{\kappa\gamma}{2(2-\kappa)}>0,$
thus $h$ is 2-supercoercive. ∎
As a consequence, we have the following result (which includes also the case $\gamma=0$). Since the proof is standard, it is omitted.
@@
Let $K\subset\mathbb{R}^{n}$ be a closed convex set and $h:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ be a proper, lsc and $(\kappa,\gamma)$-strongly quasar-convex function with respect to $\overline{x}\in{\rm argmin}_{\mathbb{R}^{n}}\,h$ ($\kappa\in\,]0,1]$ and $\gamma\geq 0$) on $K\subset{\rm dom}\,h$. Then $h(\cdot)+\frac{1}{2\beta}\lVert\cdot\rVert^{2}$ is lsc and $2$-supercoercive on $K$ for all $\beta>0$. As a consequence, ${\rm Prox}_{\beta h}(K,z)$ is nonempty and compact for all $z\in\mathbb{R}^{n}$ and all $\beta>0$.
@@
    In contrast to the convex and (strongly) quasiconvex cases (see [8, Proposition 16.44] and [30, Proposition 7], respectively), relation (24) holds only for $\overline{x}\in{\rm argmin}_{\mathbb{R}^{n}}\,h$. Surprisingly, as we will see in the next section, the information provided by (24) is enough to guarantee (linear) convergence to the optimal solution of a (strongly) quasar-convex function and to provide a detailed complexity analysis.
@@
    Recall that if $K\subset\mathbb{R}^{n}$ is a closed set and $h:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is a proper function with $K\cap{\rm dom},h\neq\emptyset$, then $h$ is said to be prox-convex on $K$ (see [15, 16]) if there exists $\alpha>0$ such that for every $z\in K$, ${\rm Prox}_{h}(K,z)\neq\emptyset$, and
@@
The following result provides a connection between fixed points of the proximity operator and minimal points of quasar-convex functions.
@@
Let $K\subset\mathbb{R}^{n}$ be a closed convex set and $h:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ be a proper, lsc and $(\kappa,\gamma)$-strongly quasar-convex function with respect to ${\rm argmin}_{K}\,h$ ($\kappa\in\,]0,1]$ and $\gamma\geq 0$) on $K\subset{\rm dom}\,h$, and $\beta>0$. Then
@@
$(\subset)$ Let $x^{*}\in{\rm Prox}_{\beta h}(K,x^{*})$. Then, $h(x^{*})\leq h(x)+\frac{1}{2\beta}\lVert x^{*}-x\rVert^{2}$ for all $x\in K$. Take $x=\lambda\overline{x}+(1-\lambda)x^{*}\in K$ with $\lambda\in[0,1]$ and $\overline{x}\in{\rm argmin}_{K}\,h$. Since $h$ is strongly quasar-convex with modulus $\kappa\in\,]0,1]$ and $\gamma\geq 0$, we have

### v2

@@ 1 Introduction
Despite its simplicity, PPA is a powerful tool in continuous optimization, as it allows us to simplify the original problem by solving better-conditioned subproblems. For instance, when $h$ is convex, the function $x\mapsto h(x)+\frac{1}{2\beta_{k}}\lVert x^{k}-x\rVert^{2}$ is strongly convex. This property makes solving the subproblems easier because proper, lower semicontinuous (lsc henceforth) strongly convex functions possess remarkable properties, such as a unique minimizer and quadratic growth. In fact, when $h$ is convex, the sequence $\{x^{k}\}_{k}$ generated by (2) is a minimizing sequence for problem (1). Specifically, $\{h(x^{k})\}_{k}$ is decreasing, converges to the minimum (though not linearly), and achieves a complexity of $\mathcal{O}(\varepsilon^{-1})$. Moreover, if $h$ is strongly convex, then $\{x^{k}\}_{k}$ converges linearly to the unique solution with an improved complexity rate of $\mathcal{O}(\ln(\varepsilon^{-1}))$.
@@
Our Contribution: We study (strongly) quasar-convex functions with modulus $\gamma\geq 0$. To that end, first we establish sufficient conditions for (nonsmooth) functions to be (strongly) quasar-convex. In particular, we show that the $\ell_{p}$-regularization ($0<p<1$) [52] and the CES (Constant Elasticity of Substitution) [3], Cobb-Douglas [12], and Leontief [33] production/utility functions are quasar-convex (and strongly quasar-convex on bounded convex sets). Second, we investigate the proximity operator of quasar-convex functions by proving that it is always nonempty (i.e., existence of iterates is guaranteed) and demonstrating cases where it is not a singleton. Furthermore, we establish a useful relationship between the minimizers of the proximity operator and the minimizers of the function. Third, we prove that the set of fixed points of the proximity operator coincides with the set of minimizers of a quasar-convex function, which provides the stopping criterion for PPA. Fourth, we implement PPA for quasar-convex functions ($\gamma\geq 0$) and show that the generated sequence is minimizing with a complexity of $\mathcal{O}(\varepsilon^{-1})$ when $\gamma=0$, matching the rate for convex functions [19]. Fifth, for the special case of strongly quasar-convex functions ($\gamma>0$), we prove that PPA converges linearly to the unique solution and achieves an $\mathcal{O}(\ln(\varepsilon^{-1}))$ complexity rate, exactly as in the strongly convex case. Finally, we present numerical experiments with nonconvex examples to demonstrate the potential of our results.
@@ 2 Preliminaries and Basic Definitions
The inner product of $\mathbb{R}^{n}$ and the Euclidean norm are denoted by $\langle\cdot,\cdot\rangle$ and $\lVert\cdot\rVert$, respectively. The open ball with center at $x_{0}\in\mathbb{R}^{n}$ and radius $\varepsilon>0$ is denoted by $\mathbb{B}(x_{0},\varepsilon)$. We denote $\mathbb{R}_{+}:=[0,+\infty[$ and $\mathbb{R}_{++}:=\,]0,+\infty[$, hence $\mathbb{R}^{n}_{+}:=[0,+\infty[\,\times\ldots\times[0,+\infty[$ and $\mathbb{R}_{++}^{n}:=\,]0,+\infty[\,\times\ldots\times\,]0,+\infty[$ ($n$ times), respectively. Let $K$ be a nonempty set in $\mathbb{R}^{n}$, its closure is denoted by $\overline{K}$ and its asymptotic (recession) cone by $K^{\infty}:=\{u\in\mathbb{R}^{n}:~\exists~t_{k}\rightarrow+\infty,~\exists~x^{k}\in K,~\frac{x^{k}}{t_{k}}\rightarrow u\}$. Furthermore, for any set $K\subset\mathbb{R}^{n}$, it follows from [5, Proposition 2.1.2] that $K^{\infty}=\{0\}$ if and only if $K$ is bounded.
@@
    (strongly) quasiconvex on ${\rm dom}\,h$, if there exists $\gamma\geq 0$ such that, for all $x,y\in{\rm dom}\,h$ and all $\lambda\in[0,1]$, we have
@@
Let $K\subset\mathbb{R}^{n}$. A proper function $h:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is said to be:
@@
    2-supercoercive on $K$, if
    $\liminf_{\begin{subarray}{c}x\in K,\,\lVert x\rVert\rightarrow+\infty\end{subarray}}\frac{h(x)}{\lVert x\rVert^{2}}>0,$ (7)
@@
    coercive on $K$, if
    $\lim_{\begin{subarray}{c}x\in K,\,\lVert x\rVert\rightarrow+\infty\end{subarray}}h(x)=+\infty.$ (8)
    or equivalently, if $S_{\lambda}(h)=\{x\in K:\ h(x)\leq\lambda\}$ is bounded for all $\lambda\in\mathbb{R}$.
@@
The following generalized convexity notion has been used in several applications from machine learning in virtue of its properties for fast convergence of gradient-type methods (see [20, 23, 24, 25, 50] and references therein).
@@
Let $K\subset\mathbb{R}^{n}$ be a closed convex set, $\kappa\in\,]0,1]$, $\gamma\geq 0$, and $h:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be a proper function such that $K\subset{\rm dom}\,h$. We say that $h$ is $(\kappa,\gamma)$-strongly quasar-convex on $K$ with respect to $\overline{x}\in{\rm argmin}_{K}\,h$ if
$h(\lambda\overline{x}+(1-\lambda)x)\leq\kappa\lambda h(\overline{x})+(1-\kappa\lambda)h(x)-\lambda\left(1-\frac{\lambda}{2-\kappa}\right)\frac{\kappa\gamma}{2}\|x-\overline{x}\|^{2},$ (9)
for all $\lambda\in[0,1]$ and all $x\in K$. Furthermore, when $\gamma=0$, we simply say that $h$ is $\kappa$-quasar-convex on $K$ with respect to $\overline{x}\in\operatorname{argmin}_{K}h$.
Note that ${\rm argmin}_{K}\,h$ is a singleton when $h$ is strongly quasar-convex on $K$ ($\gamma>0$), thus in this case we say that $h$ is strongly quasar-convex with modulus $\kappa\in\,]0,1]$ and $\gamma>0$. If $h$ is quasar-convex ($\gamma=0$), then ${\rm argmin}_{K}\,h$ is star-shaped (see [25, Observation 3]), i.e., there exists an $\overline{x}_{0}\in{\rm argmin}_{K}\,h$ such that for all $\overline{x}\in{\rm argmin}_{K}\,h$ and all $\lambda\in[0,1]$, we have $\lambda\overline{x}_{0}+(1-\lambda)\overline{x}\in{\rm argmin}_{K}\,h$. Furthermore, if $\kappa=1$, then (strongly) quasar-convex functions are (strongly) star-convex [42]. In particular, every (strongly) convex function is $(1,\gamma)$-strongly quasar-convex ($\gamma\geq 0$).
@@
where “*” denotes that ${\rm argmin}_{K\,}h\neq\emptyset$ is required. All the reverse statements do not hold in general (see [20, 24, 25, 42]).
@@
Note that if $h$ is strongly quasar-convex on $K$ with modulus $\kappa\in\,]0,1]$ and $\gamma>0$, then (see [25, Corollary 1])
$h(\overline{x})+\frac{\kappa\gamma}{2(2-\kappa)}\lVert y-\overline{x}\rVert^{2}\leq h(y),~\forall~y\in K,$ (12)
where $\overline{x}\in{\rm argmin}_{K}\,h$, thus strongly quasar-convex functions satisfy a quadratic growth condition with modulus $\frac{\kappa\gamma}{2(2-\kappa)}>0$.
@@ 3.1 Properties and Examples
We begin with two auxiliary lemmas that will be used to construct explicit examples of homogeneous (strongly) quasar-convex functions. The first analyzes auxiliary functions while the second uses them to characterize quasar-convexity and strong quasar-convexity for homogeneous functions. These results will be used in the sequel.
@@
The proof of Lemma 3 follows from a direct direct analysis of the functions $\theta_{\alpha}$ and $Q_{\alpha,\kappa}$, and is therefore omitted.
@@
    The function $h$ is $\kappa$-quasar-convex on $K$ with respect to $\overline{x}=0$ if and only if
@@
    The function $h$ is $(\kappa,\gamma)$-strongly quasar-convex on $K$ with respect to $\overline{x}=0$ if and only if (17) holds and
    $0\leq\gamma\leq Q_{\alpha,\kappa}(\lambda)\frac{h(x)}{\lVert x\rVert^{2}},~\forall~x\in K\setminus\{0\},~\forall~\lambda\in\,]0,1].$ (18)
@@
Since $\overline{x}=0\in{\rm argmin}_{K}\,h$ and $h$ satisfies (16), $h(0)=0$. Thus, $h$ is $(\kappa,\gamma)$-strongly quasar-convex with respect to $\overline{x}=0$ if and only if
@@
    If $0<\alpha<1$, then $h$ is $\kappa$-quasar-convex on $K$ with respect to $\overline{x}=0$ for every $\kappa\in\,]0,\alpha]$.
@@
    If $\alpha\geq 1$, then $h$ is $\kappa$-quasar-convex on $K$ with respect to $\overline{x}=0$ for every $\kappa\in\,]0,1]$.
@@
    If $0<\alpha\leq 1$, then $h$ is $(\kappa,\gamma)$-strongly quasar-convex on $K$ with respect to $\overline{x}=0$ for every $\kappa\in\,]0,\alpha[$ and
@@
    If $1<\alpha<2$, then $h$ is $(\kappa,\gamma)$-strongly quasar-convex on $K$ with respect to $\overline{x}=0$ for every $\kappa\in\,]0,1]$ and
@@
    If $\alpha=2$, then $h$ is $(\kappa,\gamma)$-strongly quasar-convex on $K$ with respect to $\overline{x}=0$ for every $\kappa\in\,]0,1]$ and
@@
Since $K_{c}:=\{x\in K:\lVert x\rVert=c\}$ is compact and $h$ is lsc, there exists $x^{0}\in K_{c}$ such that
$h(x^{0})=\min_{x\in K_{c}}h(x)=S_{c}.$
As $x^{0}\neq 0$ and $\mathrm{argmin}_{K}\,h=\{0\}$, it follows that $S_{c}>0$. Let $x\in K\setminus\{0\}$ and take $v:=\frac{cx}{\lVert x\rVert}$. By assumption (20), we have $v\in K$ whenever $\lVert x\rVert\leq c$, thus the convexity of $K$ ensures $v\in K$ when $\lVert x\rVert>c$. Then we have two cases:
@@
$h_{p}(x)=\lVert x\rVert_{p}^{p}=\sum^{n}_{i=1}\lvert x_{i}\rvert^{p}.$ (21)
@@
It follows from Proposition 5 that $h_{p}$ is quasar-convex on $\mathbb{R}^{n}$ with $\kappa\in\,]0,p]$. Moreover, using the notation of Proposition 6, and restricting $h_{p}$ to $K=\overline{\mathbb{B}}(0,1)$ with $c=1$, we obtain $S_{1}=\min_{\|x\|=1}\|x\|_{p}^{p}=1.$ Therefore, Proposition 6 ensures that $h_{p}$ is strongly quasar-convex on $\overline{\mathbb{B}}(0,1)$ with
@@ 3.2 On the Proximity Operator
The next result provides an analysis of the behaviour at infinity of strongly quasar-convex functions.
@@
Let $K\subset\mathbb{R}^{n}$ be a closed convex set and $h:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ be a proper lsc function such that $K\subset{\rm dom}\,h$. If $h$ is a strongly quasar-convex function on $K$ with modulus $\kappa\in\,]0,1]$ and $\gamma>0$, then $h$ is $2$-supercoercive on $K$.
@@
Let $\{x^{k}\}_{k}\subset K$ be such that $\lVert x^{k}\rVert\rightarrow+\infty$. Then, we may assume that $\frac{x^{k}}{\lVert x^{k}\rVert}\rightarrow u\in K^{\infty}$ with $\lVert u\rVert=1$. Since $h$ has a minimum $\overline{x}\in K$, $h(\overline{x})\leq h(x^{k})$ for all $k\in\mathbb{N}$. Using relation (12), we have
$h(x^{k})\geq h(\overline{x})+\frac{\kappa\gamma}{2(2-\kappa)}\lVert x^{k}-\overline{x}\rVert^{2}$ $\Longrightarrow\frac{h(x^{k})}{\lVert x^{k}\rVert^{2}}\geq\frac{h(\overline{x})}{\lVert x^{k}\rVert^{2}}+\frac{\kappa\gamma}{2(2-\kappa)}\left\|\frac{x^{k}}{\lVert x^{k}\rVert}-\frac{\overline{x}}{\lVert x^{k}\rVert}\right\|^{2}$
$\Longrightarrow\liminf_{k\rightarrow+\infty}\frac{h(x^{k})}{\lVert x^{k}\rVert^{2}}\geq\frac{\kappa\gamma}{2(2-\kappa)}>0,$
thus $h$ is 2-supercoercive on $K$. ∎
As a consequence, we have the following result (which includes also the case $\gamma=0$).
@@
Let $K\subset\mathbb{R}^{n}$ be a closed convex set and $h:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ be a proper, lsc and $(\kappa,\gamma)$-strongly quasar-convex function with respect to $\overline{x}\in{\rm argmin}_{K}\,h$ ($\kappa\in\,]0,1]$ and $\gamma\geq 0$) on $K\subset{\rm dom}\,h$. Then $h(\cdot)+\frac{1}{2\beta}\lVert\cdot\rVert^{2}$ is lsc and $2$-supercoercive on $K$ for all $\beta>0$. As a consequence, ${\rm Prox}_{\beta h}(K,z)$ is nonempty and compact for all $z\in\mathbb{R}^{n}$ and all $\beta>0$.
Proof.
Since $h$ has a minimum on $K$, the sum $h(\cdot)+\frac{1}{2\beta}\lVert\cdot\rVert^{2}$ is lsc and $2$-supercoercive on $K$ for all $\beta>0$. Hence, ${\rm Prox}_{\beta h}(K,z)$ is nonempty and compact for all $z\in\mathbb{R}^{n}$ and all $\beta>0$ as a consequence of the Weierstrass Theorem. ∎
@@
    In contrast to the convex and (strongly) quasiconvex cases (see [8, Proposition 16.44] and [30, Proposition 7], respectively), relation (25) holds only for $\overline{x}\in{\rm argmin}_{K}\,h$. Surprisingly, as we will see in the next section, the information provided by (25) is enough to guarantee (linear) convergence to the optimal solution of a (strongly) quasar-convex function and to provide a detailed complexity analysis.
@@
    Recall that if $K\subset\mathbb{R}^{n}$ is a closed set and $h:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is a proper function with $K\cap{\rm dom}\,h\neq\emptyset$, then $h$ is said to be prox-convex on $K$ (see [15, 16]) if there exists $\alpha>0$ such that for every $z\in K$, ${\rm Prox}_{h}(K,z)\neq\emptyset$, and
@@
  • $(iii)$
    Observe that the function in Example 8 is strongly quasar-convex on bounded convex sets while being neither weakly convex, nor DC, nor strongly quasiconvex, i.e., the results from [2, 4, 7, 13, 26, 30, 44, 47] and references therein cannot be applied to this function.
The following result provides a connection between fixed points of the proximity operator and minimal points of quasar-convex functions. We emphasize that function $h$ does not need to be quasar-convex with respect to every point in ${\rm argmin}_{K}\,h$, only with respect to one point $\overline{x}\in{\rm argmin}_{K}\,h$ is enough.
@@
Let $K\subset\mathbb{R}^{n}$ be a closed convex set and $h:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ be a proper, lsc and $(\kappa,\gamma)$-strongly quasar-convex function with respect to $\overline{x}\in{\rm argmin}_{K}\,h$ ($\kappa\in\,]0,1]$ and $\gamma\geq 0$) on $K\subset{\rm dom}\,h$, and $\beta>0$. Then
@@
$(\subset)$ Let $x^{*}\in{\rm Prox}_{\beta h}(K,x^{*})$. Then, $h(x^{*})\leq h(x)+\frac{1}{2\beta}\lVert x^{*}-x\rVert^{2}$ for all $x\in K$. Take $x=\lambda\overline{x}+(1-\lambda)x^{*}\in K$ with $\lambda\in[0,1]$ and $\overline{x}\in{\rm argmin}_{K}\,h$. Since $h$ is $(\kappa,\gamma)$-strongly quasar-convex on $K$ with respect to $\overline{x}$, we have